OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? Please answer to and show me how you got it?

OpenStudy (anonymous):

Wasn't this asked 2 hours ago? http://openstudy.com/users/zakiaboo#/updates/4fe3872ee4b06e92b8725368

OpenStudy (anonymous):

D+Q=46 OR D=46-Q .10D+.25Q=7 .10(46-Q)+.25Q=7 4.6-.10Q+.25Q=7 .15Q=7-4.6 .15Q=2.4 Q=2.4/.15 Q=16 QUARTERS. D=46-16=30 DIMES. PROOF: .10*30+.25*16=7 3+4=7 7=7

OpenStudy (anonymous):

yes but i didnt have time to answer it and i was still confused

OpenStudy (anonymous):

Thank you! Sheaangel ^_^